### v1

@@ 4.1 Problem Definition
Our objective is to develop an online planning policy model capable of generating a time-optimal quadrotor trajectory that connects updated waypoints. As shown in Figure 1, we consider the scenario in which the vehicle follows a trajectory $p(t)=\begin{bmatrix}{{}p_{r}(t)}^{\intercal}&p_{\psi}(t)\end{bmatrix}^{\intercal}$ that traverses the $m+1$ prescribed waypoints $\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}^{0}&\cdots&\tilde{p}^{m}\end{bmatrix}$, and the waypoints are updated as $\tilde{\mathbf{p}}_{new}=\begin{bmatrix}\tilde{p}^{0}&\cdots&\tilde{p}^{k}\quad\tilde{p}_{new}^{k+1}\cdots\tilde{p}^{m}_{new}\end{bmatrix}$. We assume that the vehicle is located between $k-1$-th and $k$-th waypoints, and the waypoints are updated from the $k+1$-th waypoint onwards. The time-optimal online planning problem is then formulated to minimize the traversing time of the updated trajectory $p_{new}(t)$ as follows:
@@
The snap minimization method can be used to convert the quadrotor trajectory generation problem into a finite-dimensional optimization formulation Mellinger and Kumar (2011); Richter et al. (2016); Ryou et al. (2022). In our work, we employ the extended minimum snap trajectory formulation Ryou et al. (2022), which includes smoothness weights between polynomial pieces in the time allocation parameterization. We define $\chi(\mathbf{x},\mathcal{D\,}p(T_{k}),\tilde{\mathbf{p}}_{new})$ as a function that maps time allocation and smoothness weights to the polynomial trajectory, which is obtained by solving the following quadratic programming problem:
@@
Our goal is to find the optimal time allocations and smoothness weights that satisfy the feasibility constraints and minimize the entire trajectory time. Unlike the snap minimization method (2), which avoids modeling the feasibility constraints by minimizing the same smoothness objective, the proposed algorithm takes into account the feasibility constraints and prioritizes the trajectory time minimization. The unknown feasibility constraint $p\in\mathcal{P}(\mathcal{D}p)$ is addressed through a multi-fidelity approach, with the optimization goal being to satisfy the constraint at the highest fidelity level.
Figure 1(a) provides an overview of the proposed multi-fidelity reinforcement learning procedure. During each training iteration, the planning policy generates time allocations and smoothness weights for batches of waypoint sequences. The policy is trained with reinforcement learning to minimize trajectory time while adhering to the predicted feasibility provided by the reward estimator. Among the training sequences, policy outputs with high uncertainty in feasibility prediction are selectively evaluated at their respective fidelity levels. Subsequently, the evaluation results are then incorporated into the training dataset to update the feasibility boundary model in the reward estimator. The training procedure is performed separately for each fidelity level, with varying sample sizes; fewer samples are selected for higher-fidelity models, which require more computation.
Figure 1(b) illustrates the architecture of the planning policy and the reward model. The sequence-to-sequence model is used for the planning policy, with dedicated a multi-layer perceptron module (MLP) for each fidelity level to capture the fidelity gap. In the reward estimator, all modules share parameters across fidelity levels, and compensation for the imbalanced evaluation dataset size is achieved through a multi-fidelity Gaussian process kernel that models correlations between the fidelity levels. The planning policy, along with the dedicated MLP module for the highest-fidelity level, is then deployed for real-time replanning.
@@ 4.2 Multi-Fidelity Evaluations
Figure 3: Comparison of the speed profile and reference motor commands over trajectory time duration. The leftmost figure visualizes a minimum snap trajectory (MinSnap Traj.), which maintains its shape with uniform scaling of time allocation. The second figure illustrates how reducing time allocations uniformly increases the speed profile. The figures on the right side compare the reference commands of the quadrotor’s four motors from a minimum snap trajectory with the same waypoints but different time allocation scaling. As time allocations reduce, the trajectory’s maximum and minimum motor speeds increasingly diverge from the hovering motor speed (Stationary).
@@
Figure 4: Variations in maximum and minimum motor speeds with time allocation scaling across different fidelity levels are depicted. The lowest fidelity level uses reference motor speed commands based on the ideal quadrotor dynamics model (Ideal dynamics). The second level shows simulated motor speeds (Simulation), while the third level presents speeds obtained from a series of flight experiments (Real-world flight). Lowering time allocations results in the trajectory’s maximum and minimum motor speeds diverging from the stationary hovering motor speed (Stationary). At low speeds, all fidelity levels exhibit similar motor speeds. However, as time allocations are reduced, the fidelity gap widens, leading to noticeable differences in the feasibility boundaries across the fidelity levels.
This approach utilizes three types of feasibility models: (i) an analytical feasibility check based on reference motor speeds derived from the simplified quadrotor dynamics differential flatness transform, (ii) numerical evaluations of trajectory-tracking accuracy in a flight dynamics simulation and (iii) real-world experiments. In more detail, at the lowest fidelity level, a feasible set, $\mathcal{P}_{1}(\mathcal{D}p)$, is defined as the set of all trajectories that start from the state $\mathcal{D\,}p$ and have reference motor commands that fall within the admissible range. In the second and third fidelity levels, the feasible set $\mathcal{P}_{2}(\mathcal{D\,}p)$ and $\mathcal{P}_{3}(\mathcal{D\,}p)$ include all trajectories that can be executed using the provided flight controller with a tracking error below a certain threshold. The feasibility bound can be estimated more precisely as the fidelity level increases, but the computation time grows exponentially. Consequently, it is impractical to directly employ the high-fidelity models in reinforcement learning, which demands millions of training samples. To address this issue, our method utilizes MFBO to efficiently construct a high-fidelity model based on the low-fidelity feasibility model.
@@
For slow trajectories, the differences between the three fidelity levels are minimal, but as speed increases, distinct behaviors emerge at each fidelity level. Figure 4 illustrates the variation in maximum and minimum motor speeds for the trajectory shown in Figure 3 across different fidelity levels. The maximum and minimum motor speeds are critical determinants of a trajectory’s feasibility because tracking errors increase when motor speed commands reach their boundaries and begin to saturate. As the trajectory time duration is reduced using the line search procedure in (5), the maximum and minimum motor speeds increasingly diverge from the stationary motor speed, which is the speed when the vehicle maintains hovering. At lower speeds, motor speeds are similar across all fidelity levels, but as speed increases, the motor speed at the lowest fidelity level — the reference motor speed commands — quickly diverges, whereas the motor speeds in simulations and flight experiments do so more gradually due to controllers filtering out speed command spikes. Consequently, in the trajectory depicted in Figure 3, the vehicle is capable of achieving higher speeds in the simulated environment compared to that estimated using the ideal dynamics model. It is noteworthy, however, that there may be opposite scenarios where the trajectory is slower in the simulated evaluation due to the control logic. Similarly, the discrepancies between simulated and real-world dynamics models cause the motor speeds to diverge differently between simulation and real-world scenarios. As trajectories’s speed increases, the disparity between fidelity levels becomes more pronounced. Optimal solutions for time-optimal trajectories are typically identified in areas where this fidelity gap is significant, highlighting the need for adaptation across different fidelity levels.
@@ 4.3 Dataset Generation
The training data comprises waypoint sequences commonly used in quadrotor trajectory planning, obtained by randomly sampling waypoints within topological constraints. Initially, a dataset of waypoint sequences $\mathcal{D}_{init}$ is created by randomly sampling sequences within the unit cube $[-0.5,0.5]^{3}$. These samples are then accepted if they satisfy two topological criteria, the total Menger curvature, Menger (1930), and the total distance between waypoints, as follows:
@@
Once the time allocation ratios are determined, a line-search is performed to find the minimal trajectory time $T^{MS}_{l}$ that satisfies the feasibility constraints for each fidelity level $l\in\{1,2,3\}$. Due to the computational expense of calculating minimal times for all waypoint sequences in the second and third fidelity levels, a subset of sequences is used to estimate the inter-fidelity ratio, which then approximates minimal times for the higher levels. To be specific, for the lowest-fidelity level of dataset, the trajectory times are determined for the entire dataset $\mathcal{D}_{init}$, and for the rest of the fidelity level, the trajectory times for the subset of the preceding fidelity dataset are determined ($\mathcal{D}_{3}\subset\mathcal{D}_{2}\subset\mathcal{D}_{1}=\mathcal{D}_{init})$. By comparing the trajectory times in the each fidelity dataset to the same waypoint sequence in the lowest fidelity dataset, we estimate the trajectory time ratio:
@@
where $N_{l}$ is the size of $l$-th fidelity subset $\mathcal{D}_{l}$. For the lowest-fidelity level, the time between the waypoints are obtained by multiplying the trajectory time with the time allocation ratio: $\mathbf{x}^{\text{MS}}_{1}=T^{\text{MS}}_{1}\tilde{\mathbf{x}}^{\text{MS}}$, and for the rest of the fidelity level, the time between the waypoints are approximated as $\mathbf{x}^{\text{MS}}_{l}=\alpha_{l/1}T^{\text{MS}}_{1}\tilde{\mathbf{x}}^{\text{MS}}$. The planning policy is trained to reconstruct the trajectory starting from a randomly selected midpoint of a training waypoint sequence, while the reward estimator is trained to predict the feasibility of the reconstructed trajectory.
@@ 4.4 Training Reward Estimator
where $\chi(\mathbf{x},\mathcal{D\,}p,\tilde{\mathbf{p}})$ refers to the polynomial trajectory generated with (13) and $\mathcal{P}_{l}$ is the set of feasible trajectories at the $l$-th fidelity level.
@@
To efficiently capture the correlation between different fidelity levels from a sparse dataset, the feasibility probability is modeled using a Gated Recurrent Unit (GRU) and a Multi-Fidelity Gaussian Process Classifier (MFGPC), as illustrated in Figure 1(b). The GRU is used to extract a feature vector from the given waypoints sequence $\tilde{\mathbf{p}}$, time allocation $\mathbf{x}_{t}$, and smoothness weights $\mathbf{x}_{w}$. The initial hidden state of GRU is obtained from the trajectory state $\mathcal{D\,}p$, and the last hidden output of GRU is converted into the feature vector through a multi-layer perceptron (MLP). Subsequently, MFGPC serves as the final stage in the feasibility estimation pipeline, utilizing this feature vector to predict feasibility. The parameters of GRU and MFGPC are trained through the minimization of a variational lower bound, as detailed in Equation 11.
The reward estimator is pretrained using a dataset with fictitious labels. We randomly select half of the $l$-th fidelity dataset $\mathcal{D}_{l}$ and reduce its trajectory time $T^{\text{MS}}_{l}$ by 10 percent. Although the data points with the reduced trajectory time may be feasible, they are currently labeled as infeasible, while the remaining data points are labeled as feasible. The slower trajectory is generally more likely to be feasible, so the reward estimator can approximately learn the feasibility boundary from the generated datasets. The training waypoint sequences are sliced from randomly selected midpoints and inputted to the reward estimator.
@@
During the RL training, the reward estimator chooses a subset of the dataset along with the corresponding policy outputs and annotates them with multi-fidelity evaluations. To be specific, policy outputs with a high level of uncertainty are selected, where the uncertainty is quantified using variation ratios Freeman (1965):
$u_{l}(\mathbf{x}_{t},\mathbf{x}_{w})=1-{\max\nolimits}_{c\in\{0,1\}}P_{l}(y=c|\mathbf{x},\mathcal{D}p,\tilde{\mathbf{p}}).$ (19)
@@
The number of selected samples varies between the fidelity levels, with more samples chosen for low-fidelity evaluations and only a few for high-fidelity evaluation due to the associated cost. MFGPC compensates for the imbalanced sample size by modeling the correlation between the different fidelity levels. After evaluations are completed, the feasibility model is updated using a training dataset that includes the newly assessed data points and randomly selected prior evaluations. At the first iteration of the RL training, to enhance training stability, policy outputs closest to the minimum-snap time allocation are chosen instead of those with high uncertainty.

### v2

@@ 4.1 Problem Definition
Figure 3: Overview of the proposed multi-fidelity reinforcement learning (MFRL) method: (1) Planning policy generates actions (time allocation and smoothness weights) for waypoint sequences. (2) Reward estimator predicts trajectory feasibility and estimates rewards, where reward is the product of feasibility prediction and time reduction achieved. (3) After updating the planning policy, portions of the training batch with high uncertainty in feasibility prediction are selected for further multi-fidelity evaluation. (4) Evaluation samples vary by computational cost of each fidelity level; results update the feasibility prediction model. (5) Iterative updates train a policy maximizing time reduction while maintaining feasibility. The method incorporates real-world experiments for direct deployment in real-world online planning applications.
Our objective is to develop an online planning policy model capable of generating a time-optimal quadrotor trajectory that connects updated waypoints. As shown in Figure 1, we consider the scenario in which the vehicle follows a trajectory $p(t)=\begin{bmatrix}{{}p_{r}(t)}^{\intercal}&p_{\psi}(t)\end{bmatrix}^{\intercal}$ that traverses the $m+1$ prescribed waypoints $\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}^{0}&\cdots&\tilde{p}^{m}\end{bmatrix}$, and the waypoints are updated as $\tilde{\mathbf{p}}_{new}=\begin{bmatrix}\tilde{p}^{0}&\cdots&\tilde{p}^{k}\quad\tilde{p}_{new}^{k+1}\cdots\tilde{p}^{m}_{new}\end{bmatrix}$. We assume that the vehicle is located between k1k-1-th and $k$-th waypoints, and the waypoints are updated from the $k+1$-th waypoint onwards. The time-optimal online planning problem is then formulated to minimize the traversing time of the updated trajectory $p_{new}(t)$ as follows:
@@
The snap minimization method can be used to convert the quadrotor trajectory generation problem into a finite-dimensional optimization formulation Mellinger and Kumar (2011); Richter et al. (2016); Ryou et al. (2022). In our work, we employ the extended minimum snap trajectory formulation Ryou et al. (2022), which includes smoothness weights between polynomial pieces in the time allocation parameterization. The smoothness weights selectively relax snap minimization objective between trajectory segments, increasing the representability of the trajectory and enabling more aggressive maneuvers. We define $\chi(\mathbf{x},\mathcal{D\,}p(T_{k}),\tilde{\mathbf{p}}_{new})$ as a function that maps time allocation and smoothness weights to the polynomial trajectory, which is obtained by solving the following quadratic programming problem:
@@
Our goal is to minimize trajectory time while ensuring feasibility constraints are satisfied. Specifically, trajectory feasibility is defined as maintaining tracking errors within specified thresholds during real-world flight. Unlike the snap minimization method (2) that rely solely on smoothness objectives to indirectly address feasibility, the proposed method explicitly models unknown feasibility constraints $p\in\mathcal{P}(\mathcal{D}p)$ to enhance time-optimality. With this constraints model, we train a planning policy that generates optimal time allocations and smoothness weights for any given waypoint sequence.
The algorithm integrates two complementary components: BO and RL. The BO component cost-effectively creates the training dataset for modeling feasibility constraints by identifying trajectories near decision boundaries and evaluating feasibility through real-world flight tests. To further improve efficiency, we employ MFGPC as a surrogate model, which incorporates low-fidelity evaluation results to accurately predict constraints with minimal real-world evaluations.
The RL component trains the planning policy to generate time-optimal parameters while satisfying the modeled feasibility constraints. It uses the expected trajectory time reduction as the reward function, calculated using the feasibility constraints model developed through BO. Rather than functioning as separate stages, these components function synergistically—BO leverages planning policy outputs instead of random trajectories to efficiently generate the training dataset, while RL continuously refines the policy based on the evolving constraint model. Figure 3 illustrates an overview of the proposed multi-fidelity reinforcement learning procedure.
@@ 4.2 Multi-Fidelity Evaluations
Figure 4: Comparison of the speed profile and reference motor commands over trajectory time. The leftmost figure visualizes a minimum snap trajectory (MinSnap Traj.), which maintains its shape with uniform scaling of time allocation. The second figure illustrates how reducing time allocations uniformly increases the speed profile. The figures on the right side compare the reference commands of the quadrotor’s four motors from a minimum snap trajectory with the same waypoints but different time allocation scaling. As time allocations reduce, the trajectory’s maximum and minimum motor speeds increasingly diverge from the hovering motor speed (Stationary).
@@
Figure 5: Variations in maximum and minimum motor speeds with time allocation scaling across different fidelity levels: Ideal dynamics, Simulation and Real-world flight. Lowering time allocations results in the trajectory’s maximum and minimum motor speeds diverging from the stationary hovering motor speed (Stationary). At low speeds, all fidelity levels exhibit similar motor speeds. However, as time allocations are reduced, the fidelity gap widens, leading to noticeable differences in the feasibility boundaries across the fidelity levels. Simulation and real-world flight cases are plotted until infeasibility or excessive tracking error occurs. The ideal-dynamics model is plotted further, becoming infeasible around 7.8 seconds when motor commands saturate.
We utilize three different fidelity levels of feasibility evaluation across the paper, and our optimization goal is to satisfy the feasibility constraints at the highest fidelity level. The trajectory feasibility of each fidelity level is defined as inclusion in a corresponding feasible trajectory set, defined as $\mathcal{P}_{1}$, $\mathcal{P}_{2}$, and $\mathcal{P}_{3}$ respectively. The first fidelity evaluations are based on differential flatness of the quadrotor dynamics, which enables us to transform a trajectory and its time derivatives from the output space, i.e. , position and yaw angle with derivatives, to the state and control input space. The resulting reference control input trajectory $u(t)=\zeta(p,t)$ would enable a hypothetical perfect quadrotor aircraft to track the trajectory $p$. The set of feasible trajectories at this fidelity level is defined as
$\mathcal{P}_{1}=\>\Big{\{}p\Big{|}\zeta(p,t)\in\big{[}\underline{u},\bar{u}\big{]}^{4}\;\;\;\forall t\in\left[0,T\right]\Big{\}},$ (15)
where $\underline{u}$ is the minimum, $\bar{u}$ is the maximum motor speed, and $T$ is the total trajectory time. The second fidelity evaluations are obtained using numerical simulation. The feasible set is defined as
$\mathcal{P}_{2}=\Big{\{}p\Big{|}$ $\left\lVert p_{r}(t)-r(t)\right\rVert\leq\bar{r}\;\land\;$
$|{p_{\psi}(t)}-{\psi(t)}|\leq\bar{\psi}\ \;\;\;\forall t\in\left[0,T\right]\Big{\}},$ (16)
where $\bar{r}$ is the maximum allowable Euclidean position tracking error, and $\bar{\psi}$ is the maximum allowable yaw tracking error. $r(t)$ and $\psi(t)$ are the simulated 3D position and yaw angle, respectively. At this fidelity level, FlightGoggles Guerra et al. (2019), an open-source multicopter simulator, is used with trajectory tracking control Tal and Karaman (2020) to consider stochastic noise, motor dynamics, and aerodynamic effects. It’s worth mentioning that any vehicle dynamics and IMU simulation could be used, as our proposed algorithm is not dependent on the specific simulation or evaluation method. Tracking error calculation requires non-parallelizable dynamics updates and controller simulations with denser time steps, significantly increasing computational time compared to the first fidelity level. The third fidelity evaluations are obtained from real-world experiments using a quadrotor aircraft and motion capture system. For these actual flight tests, each evaluation incorporates dynamics of the full system, including actuation and sensor systems, vehicle vibrations, unsteady aerodynamics, battery performance, and estimation and control algorithms. We use the same controller as in simulation, and perform multiple flights to account for stochastic effects. The feasible set $\mathcal{P}_{3}$ is defined identically to (16), but the 3D position $r_{\text{mocap}}(t)$ and yaw angle $\psi_{\text{mocap}}(t)$ are measured using the motion capture system:
$\mathcal{P}_{3}=\Big{\{}p\Big{|}$ $\left\lVert p_{r}(t)-r_{\text{mocap}}(t)\right\rVert\leq\bar{r}\;\land\;$
$|{p_{\psi}(t)}-{\psi_{\text{mocap}}(t)}|\leq\bar{\psi}\ \;\;\;\forall t\in\left[0,T\right]\Big{\}},$ (17)
The feasibility bound can be estimated more precisely as the fidelity level increases, but the evaluation time grows exponentially. Empirically, evaluations at the first fidelity level take about 10 milliseconds, while the second level takes roughly 2 seconds, and the third level around 2 minutes, with times changing based on trajectory lengths. Consequently, it is impractical to directly employ the high-fidelity models in RL, which demands millions of training samples. To address this issue, our method utilizes MFBO to efficiently construct a high-fidelity model based on the low-fidelity feasibility model.
@@
For slow trajectories, the differences between the three fidelity levels are minimal, but as speed increases, distinct behaviors emerge at each fidelity level. Figure 5 illustrates the variation in maximum and minimum motor speeds for the trajectory shown in Figure 4 across different fidelity levels. The maximum and minimum motor speeds are critical determinants of a trajectory’s feasibility because tracking errors increase when motor speed commands reach their boundaries and begin to saturate. As the trajectory time is reduced using the line search procedure in (5), the maximum and minimum motor speeds increasingly diverge from the stationary motor speed, which is the speed when the vehicle maintains hovering. At lower speeds, motor speeds are similar across all fidelity levels, but as speed increases, the motor speed at the lowest fidelity level — the reference motor speed commands — quickly diverges, whereas the motor speeds in simulations and flight experiments do so more gradually due to controllers filtering out speed command spikes. Consequently, in the trajectory depicted in Figure 4, the vehicle is capable of achieving higher speeds in the simulated environment compared to that estimated using the ideal dynamics model. It is noteworthy, however, that there may be opposite scenarios where the trajectory is slower in the simulated evaluation due to the control logic. Similarly, the discrepancies between simulated and real-world dynamics models cause the motor speeds to diverge differently between simulation and real-world scenarios. As trajectories’s speed increases, the disparity between fidelity levels becomes more pronounced. Optimal solutions for time-optimal trajectories are typically identified in areas where this fidelity gap is significant, highlighting the need for adaptation across different fidelity levels.
@@ 4.3 Dataset Generation
The training data comprises waypoint sequences commonly used in quadrotor trajectory planning, obtained by randomly sampling waypoints within topological constraints. Initially, a dataset of waypoint sequences $\mathcal{D}_{init}$ is created by randomly sampling sequences within the unit cube $0.5,0.5]^{3}$. These samples are then accepted if they satisfy two topological criteria, the total Menger curvature, Menger (1930), and the total distance between waypoints, as follows:
@@
Figure 6: Line search across fidelity levels. (Left) First level ($l=1$): Reduce trajectory time to motor command boundary, $\underline{u},\bar{u}]$. (Right) Second and third levels ($l=2,3$): Reduce time to tracking error threshold, $\bar{r},\bar{\psi}$.
Once the time allocation ratios are determined, a line-search, as defined in (5), is performed to find the minimal trajectory time $T^{MS}_{l}$ for each fidelity level $l\in\{1,2,3\}$ that satisfies the feasibility constraints. Figure 6 illustrates the line search procedure in each fidelity level. Due to the computational expense of calculating minimal times for all waypoint sequences in the second and third fidelity levels, a subset of sequences is used to estimate the inter-fidelity ratio, which then approximates minimal times for the higher levels. To be specific, for the lowest-fidelity level of dataset, the trajectory times are determined for the entire dataset $\mathcal{D}_{init}$, and for the rest of the fidelity level, the trajectory times for the subset of the preceding fidelity dataset are determined ($\mathcal{D}_{3}\subset\mathcal{D}_{2}\subset\mathcal{D}_{1}=\mathcal{D}_{init})$. By comparing the trajectory times in the each fidelity dataset to the same waypoint sequence in the lowest fidelity dataset, we estimate the trajectory time ratio:
@@
where $N_{l}$ is the size of $l$-th fidelity subset $\mathcal{D}_{l}$. For the lowest-fidelity level, the time between the waypoints are obtained by multiplying the trajectory time with the time allocation ratio: $\mathbf{x}^{\text{MS}}_{1}=T^{\text{MS}}_{1}\tilde{\mathbf{x}}^{\text{MS}}$, and for the rest of the fidelity level, the time between the waypoints are approximated as $\mathbf{x}^{\text{MS}}_{l}=\alpha_{l/1}T^{\text{MS}}_{1}\tilde{\mathbf{x}}^{\text{MS}}$. In the proposed simulation and real-world flight setup, $\alpha_{l/1}\in[0.95,1.05]$ across all fidelity levels and datasets. This approximates time allocation for higher-fidelity levels, avoiding costly line searches over the entire dataset. The parameter is used only for training; for testing, we conduct line searches to determine exact minimum time allocations. The planning policy is trained to reconstruct the trajectory starting from a randomly selected midpoint of a training waypoint sequence, while the reward estimator is trained to predict the feasibility of the reconstructed trajectory. During the testing phase, we use the time allocation $\mathbf{x}^{\text{MS}}_{l}$ from the highest fidelity level solution of (4) and (5) as the baseline. We also add position and yaw deviations to waypoints to assess policy robustness. Figure 7 shows the dataset usage in training and testing.
Figure 7: Waypoints dataset generation and usage in training and testing. Initial dataset: Randomly generated waypoints scaled to room size $L_{\text{space}}$, connected using snap minimization. Training: Planning policy reconstructs trajectories from random midpoints of waypoint sequences; reward estimator predicts trajectory feasibility. Testing: Position and yaw deviations added to waypoint sequences to evaluate policy robustness.
Figure 8: Overview of the planning policy and reward estimator architecture. (Left) The planning policy, based on a sequence-to-sequence model, determines time allocation and smoothness weights for input waypoints. The model comprises an encoder (bi-directional GRU), decoder (GRU), VAE, and attention modules. (Right) The reward estimator predicts feasibility of the policy outputs using a multi-fidelity Gaussian process classifier. The reward signal is calculated as a product of feasibility prediction and time reduction achieved by the policy output.
@@ 4.4 Training Reward Estimator
[…]
@@
To efficiently capture the correlation between different fidelity levels from a sparse dataset, the feasibility probability is modeled using a Gated Recurrent Unit (GRU) and a Multi-Fidelity Gaussian Process Classifier (MFGPC), as illustrated in Figure 8. The GRU is used to extract a feature vector from the given waypoints sequence $\tilde{\mathbf{p}}$, time allocation $\mathbf{x}_{t}$, and smoothness weights $\mathbf{x}_{w}$. The initial hidden state of GRU is obtained from the trajectory state $\mathcal{D\,}p$, and the last hidden output of GRU is converted into the feature vector through a multi-layer perceptron (MLP). Subsequently, MFGPC serves as the final stage in the feasibility estimation pipeline, utilizing this feature vector to predict feasibility. The parameters of GRU and MFGPC are trained through the minimization of a variational lower bound, as detailed in (11).
Figure 9: Trajectory labeling for reward estimator pretraining. (Left) Trajectories from snap minimization (dark blue region) are sped up to their feasibility limit; further acceleration makes them infeasible (grey region). The yellow shaded region represents the already traversed portion of the trajectory. (Right) After scaling, most trajectories become infeasible, but some remain feasible when reconstructed from random midpoints in the training data.
The feasibility prediction model is pretrained using a dataset with fictitious labels. We randomly select half of the $l$-th fidelity dataset $\mathcal{D}_{l}$ and reduce its trajectory time $T^{\text{MS}}_{l}$ by $\alpha_{\text{pretrain}}=0.9$. Although the data points with the reduced trajectory time may be feasible, as illustrated in Figure 9, they are currently labeled as infeasible, while the rest retain their feasible labels. This approach simulates infeasible data without risking actual infeasible trajectory generation, as all original trajectories are from the feasible set near the boundary. The model learns to approximate the feasibility boundary based on the principle that slower trajectories are generally more feasible. Once the model learns a brief structure of the feasibility boundary, it can be refined during the following MFRL procedure, which safely and efficiently collects a new dataset from scratch near the boundary. The time scaling factor $\alpha_{\text{pretrain}}$ balances two extremes: if too small, the model can’t differentiate near-boundary trajectories; if too large, it lacks data at speed extremes. This parameter is empirically chosen to maximize MFRL training stability, as extreme values may cause the model to converge to suboptimal solutions. For training, waypoint sequences are inputted to the model starting from randomly selected midpoints as shown in Figure 7.
@@
During the RL training, we employ BO to train the reward estimator. Specifically, the reward estimator chooses a subset of the dataset along with the corresponding policy outputs and annotates them with multi-fidelity evaluations. To be specific, policy outputs with a high level of uncertainty are selected, where the uncertainty is quantified using variation ratios Freeman (1965):
[…]
@@
The number of selected samples varies between the fidelity levels, with more samples chosen for low-fidelity evaluations and only a few for high-fidelity evaluation due to the associated cost. MFGPC compensates for the imbalanced sample size by modeling the correlation between the different fidelity levels. After evaluations are completed, the feasibility model is updated using a training dataset that includes the newly assessed data points and randomly selected prior evaluations. The subset selection and reward estimator training are performed sequentially across each fidelity level; the same data points may be evaluated at multiple fidelity levels. At the first iteration of the RL training, to enhance training stability, policy outputs closest to the minimum-snap time allocation are chosen instead of those with high uncertainty.